(6.1) Find det(C) if (1) C =  λ λ + 1 λ λ − 1  (6.2) Use the cofactor expansion to determine 2 0 0 0 3 1 2 0 2 −5 0 4 1 3 0 3 (6.3) Consider the matrix A =  1 4 2 3  (a) Compute A −1 (b) Find det(A **−1 ) (c) Deduce a relation (if it exists) between det(A) and det(A **−1 


1
Expert's answer
2021-05-11T08:32:03-0400

6.1. 


"C=\begin{bmatrix}

 \lambda& \lambda+1\\

 \lambda& \lambda-1

 \end{bmatrix}"


"Det(C)=\begin{vmatrix}

 \lambda& \lambda+1\\

 \lambda& \lambda-1

 \end{vmatrix}"


    "=\lambda(\lambda-1)-\lambda(\lambda+1)

 =-2\lambda"

6.2

        

The given determinant is "\begin{vmatrix}

 2&0&0&0\\

 3&1&2&0\\

 2&-5&0&4\\

 1&3&0&3

 \end{vmatrix}"


Use the cofactor expansion corresponding to the first row.


"=2 \begin{vmatrix}

 1&2&0\\

 -5&0&4\\

 3&0&3

 

 \end{vmatrix} -0+0-0"


"=2 \begin{vmatrix}

 1&2&0\\

 -5&0&4\\

 3&0&3

 

 \end{vmatrix} 



 =2[1(0-0)-2(-15-12)+0]

 =2[54]=108"


6.3

"(a) \text{The given matrix is-
}
A=\begin{bmatrix}

 1&4\\

2&3\end{bmatrix}"


"det A=\begin{vmatrix}

 1&4\\

2&3\end{vmatrix}=3-8=-5"


"det A=-5 \neq 0."


"\Rightarrow A^{-1}" exist.


Now, "adj.(A)=\begin{bmatrix}

 3&-2\\

2&3\end{bmatrix}^T



\Rightarrow adj(A)=\begin{bmatrix}

 3&-4\\

-2&1\end{bmatrix}"



So, "A^{-1} =\dfrac{adj A}{det (A)}=\dfrac{1}{-5}\begin{bmatrix}

 3&-4\\

-2&1\end{bmatrix}=\begin{bmatrix}

 \dfrac{-3}{5} & \dfrac{4}{5}\\\\

\dfrac{2}{5} & \dfrac{-1}{5}\end{bmatrix}"

(b) 


"det(A^{-1})=\begin{vmatrix}

 \dfrac{-3}{5} & \dfrac{4}{5}\\\\

\dfrac{2}{5} & \dfrac{-1}{5}\end{vmatrix}"


    "=\dfrac{3}{25}-\dfrac{8}{25}=\dfrac{3-8}{25}=\dfrac{-5}{25}=\dfrac{-1}{5}"



(c) We have, "det(A)=-5 \text{ and }det(A^{-1})=\dfrac{-1}{5}"


  "det(A).det(A^{-1})=(-5).(-\dfrac{1}{5})=1"


  Hence "det(A).det(A^{-1})=1"


This is the required relation.